# Sustainability and Carbon Emissions of Future Accelerators

### v2

@@ 6.1 Negative emissions technologies
Negative emissions technologies can be separated into two broad categories: nature-based solutions and carbon dioxide capture [80]. Note that Carbon Capture and Storage (CCS) is not a negative emissions technology, as it aims to capture the CO2 directly at the source, so its contribution is carbon neutral rather than carbon negative [81]. Standard nature-based technologies for negative emissions, as discussed by the Green ILC [82] project, is afforestation (planting new trees) or re-forestation (reintroducing a forest), which absorbs CO2 as part of the photosynthesis process. Note that we do not consider protecting an already existing forest from being cut down in the future to be net negative in terms of emissions. Trees absorb CO2 according to their species and age. A juvenile tree (15 years old) will start to reach peak absorption until it reaches the mature stage (50 years old) where, depending on the species, it could absorb between 200-400 kg of CO2 per year [83]. In comparison, during the sapling stage (5-15 years old), the tree would absorb between 10-50 kg of CO2 per year. Another important aspect to consider in using afforestation is that the overall carbon cycle of a tree during its whole life span is carbon neutral, so all of the carbon absorbed during its life will be released into the atmosphere when the tree dies and decays. For trees to be a negative emissions technology, they need to be cut and used, for example in construction, so as to sequester the carbon. If the trees are burned and used as a source of fuel, then CCS needs to be utilized to capture those emissions. In order to absorb 1 million tonnes of CO2 emissions over a period of 10 years (an amount equivalent to that which would be released through the construction of the FCC-ee tunnel), we estimate that about 10 million new trees would be needed, which would cover an area between 4,000 to 10,000 hectares of land depending on various factors including the tree species. This area is equivalent to a factor 10 to 30 of the area covered by Central Park in New York City, or four to 11 times the area of Richmond Park in the London area. After those 10 years, the trees would need to be cut and used for e.g. construction to avoid the re-release of CO2.
@@
The second approach to negative emissions technologies is that of carbon capture, in particular Direct Air Capture (DAC). There are currently two main methods for DAC. One uses a hydroxide solution to capture CO2 from ambient air, which is then heated to high temperatures (using renewable power) to release the CO2 so it can be stored and the hydroxide solution reused. The other method uses amine adsorbents in small, modular reactors; the temperature needed to re-release the CO2 is estimated to be less than for the first method [84]. The latter method is used by the Swiss firm Climeworks [85]. It currently operates the Orca DAC plant in Iceland which can capture 4,000 tonnes of CO2 per year, and a larger facility, Mammoth, is currently being built which will capture four times this amount per year, with construction costs estimated at \$600M. Climeworks uses between 8,000 to 18,000 kWh of electrical energy to capture 1 tonne of CO2. In Iceland this power is provided by the Hellisheiði geothermal facility, which can produce 2.65 TWh of electricity per year. It would then take Mammoth 28 years of operation in order to absorb 1 million tonnes of CO2 emissions from the atmosphere.
@@ 7 SUMMARY AND CONCLUSION
It is likely that gas detectors will continue to be an attractive option as part of the future detectors associated with those energy-frontier colliders. Unfortunately, the search for gases with low GWP and acceptable toxicity is challenging and it is imperative the particle physics detector community fully engage in this R&D. Suitable replacement of those detectors with non-gas options should also be explored. The computing section highlighted the growing importance of the embodied carbon in computer components as power continues to become carbon free, while the travel section conveyed the importance of reducing travel and finding innovative ways to pursue collaborative research.
The central message of these findings is that significantly reducing emissions or adhering to a net zero target for these future projects (using current level of research budgets) will be extremely challenging. The physics program of all these projects demonstrate how vibrant, creative and exciting the future of the field is. In order to secure this bright future it is imperative for the particle physics community to take actions for sustainability that go beyond our traditional accelerator and detector activities. Although the techniques associated with, for example, producing low carbon cement and direct air capture have been demonstrated, their application at scale is far from guaranteed and requires technological innovation that the particle physics community can positively contribute to. Our future depends on it.

### v3

@@ 6.1 Negative emissions technologies
Negative emissions technologies can be separated into two broad categories: nature-based solutions and carbon dioxide capture [83]. Note that Carbon Capture and Storage (CCS) is not a negative emissions technology, as it aims to capture the CO2 directly at the source, so its contribution is carbon neutral rather than carbon negative [84]. Standard nature-based technologies for negative emissions, as discussed by the Green ILC [85] project, is afforestation (planting new trees) or re-forestation (reintroducing a forest), which absorbs CO2 as part of the photosynthesis process. Note that we do not consider protecting an already existing forest from being cut down in the future to be net negative in terms of emissions. Trees absorb CO2 according to their species and age. A juvenile tree (15 years old) will start to reach peak absorption until it reaches the mature stage (50 years old) where, depending on the species, it could absorb between 200-400 kg of CO2 per year [86]. In comparison, during the sapling stage (5-15 years old), the tree would absorb between 10-50 kg of CO2 per year. Another important aspect to consider in using afforestation is that the overall carbon cycle of a tree during its whole life span is carbon neutral; all of the carbon absorbed during its life will be released into the atmosphere when the tree dies and decays. For trees to be a negative emissions technology, they need to be cut and used, for example in construction, so as to sequester the carbon. If the trees are burned and used as a source of fuel, then CCS needs to be utilized to capture those emissions. In order to absorb 1 million tonnes of CO2 emissions over a period of 10 years (an amount equivalent to that which would be released through the construction of the FCC-ee tunnel), we estimate that about 10 million new trees would be needed. These would cover an area between 4,000 to 10,000 hectares of land, depending on various factors including the tree species. This area is equivalent to a factor 10 to 30 of the area covered by Central Park in New York City, or four to 11 times the area of Richmond Park in the London area.
@@
The second approach to negative emissions technologies is that of carbon capture, in particular Direct Air Capture (DAC). There are currently two main methods for DAC. One uses a hydroxide solution to capture CO2 from ambient air, which is then heated to high temperatures (using renewable power) to release the CO2 so it can be stored and the hydroxide solution reused. The other method uses amine adsorbents in small, modular reactors; the temperature needed to re-release the CO2 is estimated to be less than for the first method [87]. The latter method is used by the Swiss firm Climeworks [88]. It currently operates the Orca DAC plant in Iceland which can capture 4,000 tonnes of CO2 per year. A larger facility, Mammoth, which will capture four times this amount per year, is currently being built with construction costs estimated at \$600M. Climeworks uses between 8,000 to 18,000 kWh of electrical energy to capture 1 tonne of CO2. In Iceland this power is provided by the Hellisheiði geothermal facility, which can produce 2.65 TWh of electricity per year. It would then take Mammoth 28 years of operation in order to absorb 1 million tonnes of CO2 emissions from the atmosphere.
@@ 7 SUMMARY AND CONCLUSION
It is likely that gas detectors will continue to be an attractive option as part of the future detectors associated with those energy-frontier colliders. Unfortunately, the search for gases with low GWP and acceptable toxicity is challenging and it is imperative that the particle physics detector community fully engage in this R&D. Suitable replacement of those detectors with non-gas options should also be explored. The computing section highlighted the growing importance of the embodied carbon in computer components as power continues to become carbon free, while the travel section conveyed the importance of reducing travel and finding innovative ways to pursue collaborative research.
The central message of these findings is that significantly reducing emissions or adhering to a net zero target for these future projects (under current research budgets) will be extremely challenging. The physics program of all these projects demonstrate how vibrant, creative and exciting the future of the field is. In order to secure this bright future it is imperative for the particle physics community to take actions for sustainability that go beyond our traditional accelerator and detector activities. Although the techniques associated with, for example, producing low carbon cement and direct air capture have been demonstrated, their application at scale is far from guaranteed and requires technological innovation that the particle physics community can positively contribute to. Our future depends on it.